\begin{thebibliography}{99}

\bibitem[BB94]{bb}
Bryan Birch
\newblock Noncongruence Subgroups, Covers and Drawings.
\newblock The Grothendieck Theory of Dessins d’Enfant London Math. Soc. Lecture Notes 200, Cambridge Univ. Press, 1994

\bibitem[AG85]{ag}
A. Grothendieck, 
\newblock ``Esquisse d'un Programme", 
\newblock Preprint 1985.

\bibitem[LK08]{lk}
Chris A. Kurth  and Ling Long, 
\newblock Computations with finite index subgroups of ${\rm PSL}(2, \Z)$ using Farey symbols.
\newblock Advances in algebra and combinatorics, 225-242, World Sci. Publ., Hackensack, NJ, 2008.

\bibitem[Kul91]{kul}
R.S. Kulkarni, 
\newblock An arithmetic-geometric method in the study of subgroups of the modular group,
\newblock Amer. J. Math. 113 (1991) 1053–1134.

\bibitem[Kul83]{rk}
R. Kulkarni, 
\newblock An extension of a theorem of Kurosh and applications to Fuchsian groups, 
\newblock Michigan Math. J.(30) 1983, 259-272.

\bibitem[Kul83]{rk1}
Ravi S. Kulkarni,
\newblock A New Proof and an Extension of a Theorem of Millington on the Modular Group.
\newblock Bull. London Math. Soc. 17 (1985) 458-462.

\bibitem[LTL95]{tl} 
M.L. Lang, C.H. Lim, and S.P. Tan, 
\newblock Independent generators for congruence subgroups of Hecke groups,
\newblock Math. Z. 220 (1995) 569–594.

\bibitem[LLT95]{lt} 
M.L. Lang, C.H. Lim, S.P. Tan, 
\newblock An algorithm for determining if a subgroup of the modular group
is congruence, 
\newblock J. Lond. Math. Soc. (2) 51 (1995) 491–502.

\bibitem[LL2016]{ll} 
Cheng Lien Lang and Mong Lung Lang 
\newblock Arithmetic and geometry of the Hecke groups
\newblock Journal of Algebra 460 (2016) 392-417.

\bibitem[McM23]{cm} 
 Curtis T. McMullen, 
 \newblock Billiards, Heights, and the arithmetic of non-arithmetic groups. 
 \newblock Invent. Math. 228 (2022), no. 3, 1309-1351.

\bibitem[Mil69]{mil} 
M. H. Millington, 
\newblock Subgroups of the Classical Modular Group.
\newblock J. London Math. Soc (2), 1969, 351-357.

\bibitem[Rad29]{rd}
H. Rademacher,
\newblock Uber die Erzeugenden der Kongruenzuntergruppen der Modulgruppe, Abh. Hamburg, 7(1929), 134-148.

\bibitem[Ran77]{rr} 
Robert A. Rankin,
\newblock The  Modular Group and its Subgroups.
\newblock Ramanujan Institute Madras, Lecture Notes, 1969.

\bibitem[KT77]{kt}
Kisao Takeuchi,
\newblock Arithmetic triangle groups. 
\newblock J. Math. Soc. Japan.

\bibitem[TM18]{tm}
Valdo Tatitscheff, Yang-Hui He, John McKay,
\newblock Cusps, Congruence Groups and Monstrous Dessins
\newblock Arxiv preprint https://arxiv.org/abs/1812.11752.

\bibitem[DT75]{td}
D.J.Tingley, 
\newblock Elliptic Curves and Modular Functions. 
\newblock Oxford D.Phil thesis 1975.

\bibitem[TR23]{tr}
Devendra Tiwari and Harshavardhan Reddy,
\newblock Revisiting Kulkarni's Topological proof of Millington's Theorem,
\newblock Contemporary Math. AMS 2023.


\bibitem[EH1936]{eh}
E. Hecke, 
\newblock Uber die Bestimmung Dirichletscher Reichen durch ihre Functionalgleichungen. 
\newblock Math. Ann. 112, (1936), 664–699.

\bibitem[JS78]{js}
G. A. Jones and D. Singerman,
\newblock Theory of maps on orientable surfaces. 
\newblock Proc. London Math. Soc. (3) 37 (1978), 273–307.

\bibitem[DS70]{ds}
D. Singerman,
\newblock Subgroups of Fuchsian groups and finite permutation groups.
\newblock Bull. London Math. Soc. 2 (1970), 319–323.

\bibitem[WT78]{wt}
W. Thurston, 
\newblock Geometry and Topology of 3-Manifolds. 
\newblock Princeton University Lecture Notes, 1978

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[T1]{fontenc}
\usepackage{lmodern}
\usepackage{amssymb,amsmath,amsthm,amstext,amscd,
latexsym,graphics,graphicx,bbm,caption,mathrsfs}
\usepackage{eso-pic}
\usepackage{tikz}
\usepackage{pgfplots}
\usepackage{float}
\usetikzlibrary{tikzmark}

\usepackage{caption}
\usepackage{subfigure}

\usepackage{tikz-cd}
\usepackage{fancyhdr}
\usepackage{xy,xypic,float,pgf}

\usetikzlibrary{arrows,shapes,positioning}
\usetikzlibrary{decorations.markings}
\usetikzlibrary{calc}
\usepackage{amsfonts,amsmath,amssymb}
\usepackage[english]{babel}
\usepackage[T1]{fontenc}
\usepackage[utf8x]{inputenc}
\usepackage{hyperref}
\usepackage{verbatim}
\usepackage{wrapfig}
\usepackage{graphicx,graphics}

\makeatletter
\newcommand\xleftrightarrow[2][]{%
  \ext@arrow 9999{\longleftrightarrowfill@}{#1}{#2}}
\newcommand\longleftrightarrowfill@{%
  \arrowfill@\leftarrow\relbar\rightarrow}
\makeatother
\makeatletter
\newcommand*{\encircledd}[1]{\relax\ifmmode\mathpalette\@encircledd@math{#1}\else\@encircledd{#1}\fi}
\newcommand*{\@encircledd@math}[2]{\@encircledd{$\m@th#1#2$}}
\newcommand*{\@encircledd}[1]{%
	\tikz[baseline,anchor=base]{\node[draw,circle,outer sep=0pt,inner sep=.2ex] {#1};}}
\makeatother

\newtheorem{thm}{Theorem}[section]

\newtheorem{cor}{Corollary}

\newtheorem{dfn}{Definition}[section]

\def\N{\mathbb N}

\def\R{\mathbb R}
\def\Q{\mathbb Q}
\def\Z{\mathbb Z}
\def\H{{\bf U}}

\def\PSL{{\rm PSL}}
\def\SL{{\rm SL}}
\def\T{\mathcal T^{*}_S}
\def\G{{\Gamma}}
\def\g{{\rm H_q}}
\def\D{\Delta^{*}}
\def\d{\Delta^{*'}}
\def\dd{\Delta^{*"}}
\def\E{\mathcal E_{f, S}}

\begin{document}

\title[Hecke triangle Groups and Dessin d'enfant]{Hecke triangle Groups and Dessin d'enfant} 
\author{Devendra Tiwari}
 \address{Bhaskaracharya Pratishthan, 56/14, Erandavane, Damle Path,
Off Law College Road, Pune 411004}
\email{devendra9.dev@gmail.com}
\subjclass[2000]{Primary 20H10; Secondary 51M10}
\keywords{ Modular group, Hecke Group, Graphs, Riemann Surfaces.}
\date{\today}
\begin{abstract}
In this work we will construct bipartite graphs, famously known as Dessin d'enfant, corresponding to finite index subgroups of Hecke triangle groups $(2, q, \infty )$. Then using a results of \cite{ll} we shall show the correspondences among the special polygons, the bi-partite graph, and the tree diagram for a finite index subgroup of the Hecke triangle groups $(2, q, \infty )$.
\end{abstract}
\maketitle


\section{Introduction}

The idea of representing finite index subgroups of modular group $\G = \PSL(2, \Z)$ (or corresponding coverings) by drawings, under various names, has occurred to lots of people. The best in our knowledge, the earliest one was in a manuscript of Tom Storer, dated about 1970. The drawings for the subgroups $\Gamma_0(N)$ of the modular group occur in the thesis of Dave Tingley \cite{td}. For more literature of this topic, which is related to theory of maps on Riemann surfaces we refer to \cite{bb, js, rk, tl, tm, tr} and references therein. Importance of maps lies in the Grothendieck's program \cite{ag} that there is an intimate link between this purely combinatorial object, arithmetic geometry of curves and the absolute Galois group $G_{\Q}$ of the field of algebraic numbers.

\medskip Among these works, for our present study, we have chosen Ravi Kulkarni's work \cite{kul}, in which he applies hyperbolic geometry to a classical problem in number theory of finding fundamental domains for congruence subgroups of the classical modular groups. This leads to an algorithmic construction of fundamental domain for the action of modular subgroups on the upper half plane $\H$ which has a particularly nice shape. Namely, a \emph{special polygon} is a connected polygonal region $P$ in the upper-halfplane such that all of the vertices of $P$ will be cusps and its boundary $\partial P$ consists entirely of even and odd edges together with a side-pairing satisfying certain considtions (see next section for precise definition). This work of Kulkarni substantially solves a classical problem of Rademacher \cite{rd}. The method developed in this work is of ``Farey Symbols" based on Farey subdivisions serving as endpoints of the boundary arcs of the fundamental domain. 

\medskip In the same work \cite{kul}, Kulkarni also shows that trees and graphs serve a role in classifying finite index subgroups of the Modular group $\Gamma = \PSL(2, \Z)$. More precisely he relates the conjugacy classes of a finite index subgroups of the modular group with a combinatorial gadget, known as ``bipartite cuboid graphs", which are special dessins d'enfants. 

\begin{dfn}

    A bipartite cuboid graph is a finite connected graph drawn on a Riemann surface such that its vertices are either black or white. Moreover, every edge in the graph connects a black to a white vertex and the valency (that is, the number of edges incident to a vertex) of a black vertex is either 1 or 2, that of a white vertex is either 1 or 3. Finally, for every white vertex of valency 3, there is a prescribed cyclic order on the edges incident to it. 

\end{dfn}

\medskip It is this combinatorial aspects of Kulkarni's work, which also has topological underpinnings \cite{rk} in the study of modular group, on which we focus in this work. In particular we show that {\rm
Theorem (4.2) of \cite{kul} generalizes for Hecke group, if one replaces the bi-partite cuboid 
graphs by the graphs in which every vertex has valence 1 or $q$, we call such graph \emph{bi-partite $q$-boid graph}\footnote{This term  was coined by Ser Peow Tan during his virtual talk in the year-long program at BP, Pune in December 2021.} , see section 3 for the precise definition. The Hecke groups are an obvious generalisation of the modular group $\G$. Hecke groups are class of triangle groups $\Delta (a, b, c)$ of the signature $(2, q, \infty)$ where $q \geq 3 \in \Z$, we will denote them by $\g$. They are  Fuchsian group of the first kind with signature $(2, q, \infty$) and they are isomorphic to the free products of two finite cyclic groups of orders $2$ and $q$.

\medskip For such graphs we prove the following in section 3:

\begin{thm} The conjugacy classes of subgroups of finite index in $\g$ are in 1-1 correspondence with the isomorphism classes of bipartite $q$-boid graphs. 
\end{thm}

In section 4, we define a closely related notion $q$-boid tree diagram and deduce the correspondences among the following objects associated with a finite index subgroups of $H_q$:

\begin{enumerate}
    \item Special polygons, 
\item Bi-partite $q$-boid graphs, 
\item $q$-boid tree diagrams.

\end{enumerate}

More precisely we prove the following results in section 4 

\begin{thm} There is a finite-to-one map from isomorphism classes of $q$-boid tree diagrams and isomorphism classes of bipartite $q$-boid graphs. 
 \end{thm}

 \begin{thm} There is a finite-to-one map from special polygons onto the isomorphism classes of $q$-boid tree diagrams. 
 \end{thm}


\section{Preliminaries of Hecke Groups}

\medskip In this section we will breifly review the some basic definitions and properties of Hecke groups and fix the notation, which we will requre for the proofs in the section 3. For most part we refer the work of M.L. Lang et.al., \cite{ll}, see also \cite{cm} and references therein. 

\medskip To study Dirichlet series satisfying some functional equations, Hecke introduced these groups in \cite{eh}. In \cite{ll} C.L. Lang and M.L. Lang studied Hecke groups and produced an algorithm to find the Hecke Farey Symbol, an extension of the classical Farey Symbol.   

\medskip For $q \in \N, \; \; q \geq 3$, let $\mu_q = 2\cos (\frac{\pi}{q})$, then the Hecke triangle group $\g$ is the group generated by the following Mobius transformations:

$$ z’ = A(z) = - 1/z, \; \; \; \;  z’ = B(z) = - \frac{1}{z + \mu_q}, \; \; \; q \in \N, \; \; q \geq 3. $$

Having presentation 

$$\g = \langle A, B : A^2 = B^q = e \rangle$$

\medskip These groups are discrete if and only if $\mu \geq 2$ or
$\mu = \mu_q = 2\cos (\frac{\pi}{q})$, where $q$ is an integer greater than or equal to $3$.

\medskip Adding the reflection $R(z) = 1/z$ to the generators $A$ and  $B$ of $\g$ gives the extended (or Coxeter) Hecke group with presentation

$$\g^* = \langle R, A, B : R^2 = A^2 = B^q =  (RA)^2 = (BR)^2 = e \rangle. $$

\medskip The extended Hecke group can also be generated by $P, Q$ and $R$ where $A = RP$ and $B = QR$, with defining relations 

$$ P^2 = Q ^2 = R^2 = (RP)^2 = (QR)^2 = e.$$

\medskip Geometrically, $P, Q$ and $R$ denote reflections along the sides of a triangle $\D$ denote the $(2, q, \infty)$ triangle with vertices at $i,e^{\frac{\pi i}{q}}$  and $\infty$ and interior angles equal to $\frac{\pi}{2}, \frac{\pi}{q}$ and zero. $\D$ is a fundamental domain of the Coxeter group $\g^*$. In this sense, the extended Hecke group is considered a special type of triangle group denoted by $\g^*$. Hecke group $\g$ is index two subgroup extended Hecke group $\g^*$.

\medskip Hecke groups are special because it contains cusps, which does not exists if $a, b, c < \infty$. Recall that cusps are fixed by parabolic elements, and any parabolic element is conjugate by some matrix $A \in  \SL(2, \R)$ to $ \pm \begin{bmatrix} 1 & b \\ 0 & 1 \\ \end{bmatrix}$ for some $b \in \R$. Therefore, cusps are fixed points of elements with infinite order, which does not exists in the case of $a, b, c < \infty$ where all the generators have finite order.

\subsection{Tessellation of Hecke Group}
\medskip Let $\H$ be the union of the upper half plane and the set $\{g(\infty) : g \in \g^*\}$. The $\g^*$ translates of $\D$ form a tessellation $\mathcal{T}^{*}$ of $\H$ (endowed with the hyperbolic metric) by  $\D$. 

\medskip Under the action of $\g^*$ the translates of $i, e^{\frac{\pi i}{q}}$  and $\infty$ are called even vertices, odd vertices and cusps (free vertices) of $\mathcal{T}^{*}$ respectively. Whereas the translates of the hyperbolic line joining $i$ to $\infty$ (resp. $e^{\frac{\pi i}{q}}$ to $\infty$) are called even edges (resp. odd edges) of $\mathcal{T}^{*}$. The translates of the hyperbolic line joining $i$ to $e^{\frac{\pi i}{q}}$ are called $f$-edges of $\mathcal{T}^{*}$. The hyperbolic line $(0, \infty)$ consists of two even edges. The translates of $(0, \infty)$ are called the even lines of $\mathcal{T}^{*}$. The hyperbolic line joining x and y is denoted by $(x, y)$.

\begin{dfn}
    A tessellation $\T$ on an orientable surface $S$ is a homeomorphism with the space obtained as a union of finitely many copies $\D_i$ of $\D$ where each even edge, odd edge, $f$-edge is isometrically glued to another even edge, odd edge, $f$-edge respectively so that $S$ is locally modelled on $\H$, or $Z_2 / \H$ or $Z_q / \H$ where $\Z_2, \Z_q$ act on $\H$ by a rotation around a fixed point through an angle $\pi$ or $\frac{2\pi}{q}$ respectively. 
\end{dfn}

\subsection{Special Polygon} A convex hyperbolic polygon $P$ of $\H$ is a union of some $q$-gons and a finite number of $r_i$-clusters $(r_i|q, 1 \leq r_i < q)$. The $q$-gons and the $r_i$-clusters of $P$ are called the tiles. The tiles intersect each other (if any) at either free vertices (cusps) or even lines. A special polygon $M_{\Phi} = (P, I_{\Phi})$ of $\H$ is a convex hyperbolic polygon $P$ together with a set of side
pairings $I_X$ satisfying the rules below.

\begin{enumerate}
    \item An odd edge $e$ is always paired with an odd edge $f$ (in the same $r$-cluster) and makes an internal angle $\frac{2r\pi}{q}$ with $f$. The vertex where $e$ and $f$ meet is an odd vertex of $P$. Both $e$ and $f$ are considered as sides of $P$, and are called its odd sides.

    \item Let $e$ and $f$ be two even edges in the boundary of $P$ forming an even line. Then either 
    
    \begin{enumerate}
        \item[(i)] $e$ is paired with $f$, both $e$ and $f$ are considered as sides of $P$, and are called its even sides, the point where $e$ and $f$ meet is an even vertex of $P$, or

        \item[(ii)] $e$ and $f$ form a free side of $P$, and this free side is paired with another free side of $P$.

    \end{enumerate} 

    \item $0$ and $\infty$ are vertices of $P$.
    
\end{enumerate}

\subsection{Some Results} Here we note some results from literature, which we will need in the proof of our theorems in the next section.

\medskip Then $S$ makes a complete $2$-dimensional hyperbolic orbifold in the sense of Thurston, (see chapter 13 in \cite{wt}).

\begin{thm}({\cite{kul}})\label{ku}
    The conjugacy classes of subgroups of finite index 
in $\PSL(2, \Z)$ are in 1-1 correspondence with the isomorphism classes of bipartite cuboid graphs.
\end{thm}
\begin{thm}({\cite{ll}})\label{ll}
Let H be a subgroup of finite index of $\g$. Then H has an admissible fundamental domain
\end{thm}

\section{Bipratite $q$-boid Graph}

\begin{dfn} A bipartite $q$-boid graph is a finite graph whose vertex set is divided into two disjoint subsets $V_o$ and $V_1$, such that 
\begin{enumerate}

\item[(i)] every vertex in $V_o$ has valence 1 or 2, 
\item[(ii)] every vertex in $V_1$ has valence 1 or $q$, 
\item[(iii)] there is a prescribed cyclic order on the edges incident at each vertex of valence $q$ in $V_1$ 
\item[(iv)] every edge joins a vertex in $V_o$ with a vertex in $V_1$. 

\end{enumerate}
\end{dfn}

An isomorphism of bipartite $q$-boid graphs is an isomorphism of the underlying graphs preserving the cyclic orders on the edges of each vertex of valence $q$. 

\begin{thm} The conjugacy classes of subgroups of finite index in $\g$ are in 1-1 correspondence with the isomorphism classes of bipartite $q$-boid graphs. 
\end{thm}
\begin{proof} Let $\D$ be the hyperbolic triangle described before and $S$ an orientable surface. Consider a tessellation $\T$ on $S$, we may obtain $S$ from a special polygon, associated with a finite index subgroup $\Phi$ of $\g$, by the process described in the proof of {\rm Theorem $\ref{ll}$} (see \cite{ll}), noted in previous section,  and so it is of the form $\Phi / \H$. Since $\g$ is the full group of orientation isometries of $\H$ preserving $\mathcal{T}^{*}, (S, \T)$ determines the subgroup $\Phi$ upto conjugacy in $\g$, and so the tessellation-preserving isometry classes of spaces $(S, \T)$ are in 1-1 correspondence with the conjugacy classes of subgroups of finite index in $\g$. 

\medskip Now we will prove the 1-1 correspondence of the tessellation-preserving isometry classes of the spaces $(S, \T)$ with the isomorphism classes of the bipartite $q$-boid graphs. Given $(S, \T)$ we will consider $\E$, which is the union of the $f$-edges in $S$. We can give $\E$ a natural structure of a bipartite $q$-boid graph by taking $V_0$ resp. $V_1$, to be the set of even vertices resp. odd vertices, and the cyclic order on the edges incident at a vertex of valence $q$ being the one induced from the orientation of $S$. 

\medskip Conversely when $G$ be a bipartite $q$-boid graph, consider $\D_e, \d_{e}$, which are two sets of copies of $\D$ each indexed by the edges $e$ of $G$. Attach $\D_e$ to $\d_e$ isometrically along $e$ so that a vertex in $V_0$ in one copy of $e$ is attached to a vertex in $V_0$ in the other copy of $e$. This gives us $\dd_e$ which is isometric to a hyperbolic triangle with angles $0, 0, \frac{2\pi}{q}$. There is a canonical isometry which takes $\dd_e$ to the hyperbolic triangle with vertices $0, \infty,$ and $\frac{2\pi}{q}$. Using this isometry we can give $\dd_e$ a canonical orientation and a ``counterclockwise" orientation on its boundary edges. 

\medskip If two distinct edges $e$ and $f$ share an even vertex then attach $\dd_e$ to $\dd_f$ isometrically along the complete geodesics made up of even edges in the orientation reversing way. 

\medskip If $e_1, e_2$ share an odd vertex $v$ then there will be total $q$ number of edges $e_1, e_2, \dots e_q$ also sharing $v$. By symmetry we may suppose that the cyclic order is $e_1, e_2, \dots e_q$. Orient these edges so that they ``emanate" from $v$. In $\dd_{e_i}$ using its orientation we can then uniquely determine an odd edge $e_k$ which makes the angle $+\frac{\pi}{q}$ with $e_i$. Similarly there is a unique choice of an odd edge $l$ in $\dd_{e_j}$ which makes the angle $-\frac{\pi}{q}$ with $e_j$. We attach $\dd_{e_i}$ to $\dd_{e_j}$ so that $e_k$ is isometrically identified with $e_l$. If $e$ has a terminal even (resp. odd) vertex $v$ we identify the even (resp. odd) edges of $\dd_e$ incident at $v$ in an orientation-reversing way. 

\medskip This constructions defines a $(S, \T)$ which is canonically attached to $G$ and gives the 1-1 correspondence among the isomorphism classes of bipartite $q$-boid graphs and the tessellation-preserving isometry classes of the spaces $(S, \T)$.

\end{proof}

\begin{cor} Given an index $n$ subgroup of Hecke group $\g$, the corresponding bipartite q-boid graph has $n$ edges.
\end{cor}
\begin{proof} Since for an index $n$ subgroup $\Gamma$ of $\g$, the fundamental domain of $\Gamma$ contains $n$ copies of the fundamental domain of $\g$. From the construction of the bipartite $q$-boid graphs for $\Gamma$, each edge of the bipartite $q$-boid corresponds to a copy of the fundamental domain for $\g$.
\end{proof}

\section{$q$-Boid Tree Diagrams}

In this section we will introduce a tree diagram corresponding to the bipartite graph discuss in the previous section:

\begin{dfn} A $q$-boid tree diagram is a finite tree, denoted by $T$, with at least one edge, such that 

\begin{enumerate}

\item[(i)] All the internal vertices are of valence $q$, 

\item[(ii)] There is a prescribed cyclic order on the edges incident at each internal vertex, 

\item[(iii)] The terminal vertices are partitioned into two possibly empty subsets $R$ and $B$ where the vertices in $R$ (resp. $B$) are called red (resp. blue) vertices, 

\item[(iv)] There is an involution $\sigma$ on $R$. 
\end{enumerate}
\end{dfn}

A tree diagram, as we shall show, is an useful device for the constructions of subgroups of finite index in $\g$. 

\medskip $T$ can be embedded in the plane so that the cyclic order on the edges at each internal vertex coincides with the one induced by the orientation of the plane. Any two such embeddings are in fact isotopic to each other. 

\medskip An isomorphism of two tree diagrams is defined in the obvious way and amounts to an isotopy class of planar trees satisfying (i) above. So a tree diagram can be best represented on paper without explicitly indicating the cyclic order; the red (resp. blue) vertices are represented by a small hollow (resp. shaded) circles; and distinct red vertices related by $\sigma$ are given the same numerical label, it being understood that the unlabelled vertices are fixed by $\sigma$ and different pairs of distinct red vertices related by $\sigma$ carry different labels. 

\newblock The correspondences among the special polygons, the bipartite $q$-boid graphs, and the tree diagrams are as follows. 

\begin{thm} There is a finite-to-one map from isomorphism classes of tree diagrams and isomorphism classes of bipartite $q$-boid graphs. 
 \end{thm}

 Let $T$ be a tree diagram. Identifying $v$ with $\sigma(v)$ one obtains a graph $G$. On all edges joining two internal vertices or an internal vertex with a blue vertex introduce a new vertex of valence $2$. These new vertices and the red vertices consitute $V_0$. The vertices of valence $q$ and the blue vertices constitute $V_1$ The cyclic orders on the vertices in $V_1$ are defined by ii) above. This turns $G$ into a bipartite $q$-boid graph. 

\medskip Conversely let $G$ be a bipartite $q$-boid graph. If its cycle-rank (= the first Betti number) is $r$ we can choose $r$ vertices of valence $2$ in $V_0$ so that cutting $G$ along these vertices we obtain a tree $T$. Corresponding to these $r$ cuts we have $2r$ terminal vertices in $T$. These $2r$ vertices and the terminal vertices of valence 1 in $V_0$ constitute the red vertices, and the terminal vertices in $V_1$ constitute the blue vertices. Set up the involution $\sigma$ as fixing the terminal vertices of valence 1 in $V_0$ and interchanging the two vertices obtained by each one of the $r$ cuts. We agree not to count the remaining vertices of valence 2 in $V_0$ as vertices. Finally the cyclic order on the edges incident at the vertices of valence $q$ in $T$ is the same as that in $G$. This turns $T$ into a tree diagram. Notice that $T$ depends on the choices of the $r$ cuts. 

\medskip It is clear that we have a well-defined finite-to one map from the isomorphism classes of tree diagrams onto those of bipartite cuboid graphs. 

\begin{thm} There is a finite-to-one map from special polygons onto the isomorphism classes of tree diagrams. 
 \end{thm}

Let $P$ be a special polygon and $T$ the union of all the $f$-edges in $P$. We agree not to count the even vertices in int $P$ as vertices. The even vertices resp. odd vertices in $\partial P$ constitute the red resp. blue vertices. The involution on the red vertices is given by the side-pairing datum in $P$. Finally the cyclic order on the edges incident to the vertices of valence $q$ is induced by the orientation of $P$. This turns $T$ into a tree diagram. 

\medskip Conversely let $T$ be a tree diagram. On all edges joining two internal vertices or an internal vertex with a blue vertex introduce a new vertex of valence $2$. Equip $T$ with a metric in a standard way so that each edge has the same length equal to the length of an $f$-edge (which is equal to In $q$). $T$ must have at least one red vertex or at least one blue vertex. Suppose it has a red vertex $v$. Isometrically develop the unique edge containing $v$ onto the $f$-edge joining $i$ to $p$. Then $T$ itself develops isometrically and uniquely along the $f$-edges in $\T$ so that the cyclic orders on the edges incident at the vertices of valence $q$ in $T$ match with the ones induced by the orientation of $H$. At the image of a red vertex $v$ in this development assign the even line passing through that even vertex. These even edges are paired if the vertex $v$ is fixed by the involution $\sigma$. Otherwise this complete geodesic will be considered as a free side. It will be paired with the other free side constructed at $\sigma (v)$. Similarly at the image of a blue vertex incident to the (unique) edge say $e$ assign those two $q$-edges which make an angle $\frac{\pi}{q}$ with the image of $e$. These odd edges are paired. It is easy to see that these even sides, odd sides, and free sides together with their pairing defines a special polygon. 

\medskip It is fairly clear that we have a well-defined finite-to-one map from special polygons onto the isomorphism classes of tree diagrams.

\end{document}